OpenStudy (hellokitty17):

HELP ME PLEASE!!!! Which scatter plot best represents the function f(x) = 2x + 3? A.scatter plot having y intercept at ordered pair 0, 2 and extreme points at negative 8.6, 1 and 1.6, 7 B.scatter plot having y intercept at ordered pair 0, 4 and extreme points at negative 9.8, 3 and 1.2, C. scatter plot having y intercept at ordered pair 0, 4 and extreme points at negative 8.5, 3 and 2, 7 D.scatter plot having y intercept at ordered pair 0, 2 and extreme points at negative 6.1, 1 and 2.5, 7

OpenStudy (hellokitty17):

@saseal @IrishBoy123 @iYuko @Bacon211 @juliastevens @Hero @katt9421 @BioHazard9064 @Cookie_2046 @nevermind_justschool

OpenStudy (hellokitty17):

@sebaxtiangb

hero (hero):

@satellite73 go for it.

OpenStudy (anonymous):

lol i got no idea i always thought \(y=2x+3\) is a line, not a scatter plot it would have \((0,3)\) on the graph which none of these do guess we need to see a picture to figure out which is the best approximation

OpenStudy (hellokitty17):

@Hero could you help out

hero (hero):

Go to desmos.com and plot f(x) = 2x + 3. Then plot each possible set of points from the given choices to see which set gives the best approximation of the line.

OpenStudy (anonymous):

yea plot em out to find the best fit

OpenStudy (hellokitty17):

@jcoury

OpenStudy (anonymous):

do line of best fit, basically see which one has all of the points clumped together

OpenStudy (anonymous):

can you post a screen shot? it is very hard to read a graph from a bunch of words

oregonduck (oregonduck):

she is getting it

OpenStudy (hellokitty17):

Those are the options

hero (hero):

And now we figure out that \(f(x)\) is \(2^x + 3\).

oregonduck (oregonduck):

XD

OpenStudy (hellokitty17):

@jcoury

OpenStudy (anonymous):

\[f(0)=2^0+3=1+3=4\] giving the point \((0,4)\) on the graph

OpenStudy (anonymous):

what do you need help with exactly

OpenStudy (anonymous):

try computing \(f(1)\) to get another point

OpenStudy (anonymous):

just plug in some low numbers to get the points and see which one matches

OpenStudy (hellokitty17):

which option would it be?

OpenStudy (anonymous):

i think its A

OpenStudy (hellokitty17):

ok i will just go with that. what do you guys think?

oregonduck (oregonduck):

yeah a

OpenStudy (rds8701):

I think its A too

OpenStudy (rds8701):

Can someone help me too